OpenStudy (anonymous):

solve by addition/elimination. Systems of Linear equations. Please explain as you help so that I can understand 6x-7y=-26 6x+5y=10

OpenStudy (anonymous):

just subtract equation 2 .

OpenStudy (ajprincess):

when you use elimination method u will find that one of the two variables is eliminated and u get an equation with only one variable. that is 6x-7y=-26-(1) 6x+5y=10-(2) when u subtract (1) from (2) u will get 6x+5y-6x+7y=10+26 12y=36 y=36/12 y=3. Is that clear?

OpenStudy (anonymous):

let me look....

OpenStudy (ajprincess):

hmm k. I will wait.

OpenStudy (anonymous):

ok I have been doing this different. I got the same answer. y = 3 .... the way you did it is less confusing. Now I have another question

OpenStudy (anonymous):

why did the signs change when you did that? from 10+26 instead of -10-26

OpenStudy (ajprincess):

u get -26-10 when u subtract (1) from (2).

OpenStudy (anonymous):

After I find y im suppose to solve for x and then check my work by using one of the equations

OpenStudy (anonymous):

so it would be 6x-7(3)=-26 .... so 6x=-26+21 which then makes x=5/6 but then when I check the work it does not add up .... so would I say it is not a solution and that is the answer?

OpenStudy (ajprincess):

x=-5/6 solve for x by putting the value of y in one of the eqns nd then check if the values satisfy both the eqns.

OpenStudy (anonymous):

that's what I'm trying to do but it does not

OpenStudy (anonymous):

never mind. it does work. I did not have the 5/6 as a negative. Now it works

OpenStudy (ajprincess):